OpenStudy (anonymous):

You live near a bridge that goes over a river. The underside of the bridge is an arch that can be modeled with the function y=0.000371x^2+827x, where x and y are measured in feet. How high above the river is the bridge (the top of the arch)? How long is the section of bridge above the arch?

OpenStudy (anonymous):

First question is same as finding max y, so set dy/dx = 0= 2(0.000371) x + 827 the length is how far x is to get from y = 0 at left to y = 0 at right. Solve y=0 for x and take the difference in the x values.

OpenStudy (anonymous):

I don't know how.

OpenStudy (anonymous):

if 0 = m x + b, then x = -b/m, so that's how to solve the first equation. ff y = c x^2 + e x = x( cx +e), then y = 0 at x = 0 and x+e = 0 or x = -e, so that solves the second equation, showing the length is -e {match my letters to your numbers}

OpenStudy (anonymous):

Okay...What next? Sorry, I'm so bad at math.